3. Suppose the below reaction in the forward direction is first order in A and the rate constant is 1.90×10-2 s–1. A ⇌ B. The reverse reaction is first order in B and the rate constant is 4.30×10-2 s–1 at the same temperature.
What is the value of the equilibrium constant for the below reaction at this temperature?
A ⇌ B
1
Expert's answer
2020-04-02T10:26:49-0400

Kf=1.9*10-2 s-1 forward reaction

Kb=4.30×10-2 s-1 reverse reaction

A ⇌ B

"Keq=\frac{Kf}{Kb}"

Keq=(1.9*10-2)\divide 4.30×10-2=0.44

Keq=0.44